OpenStudy (anonymous):

Fan and Medal to best answer!!!! A can contains 14 pennies, 12 nickels, 16 dimes, and 8 quarters. What is the probability of drawing either a dime or a quarter on one draw?

OpenStudy (anonymous):

A. 12/25 B. 3/5 C. 4/5 D. 12/13

OpenStudy (anonymous):

Well there are a total of 14+12+16+8 = 50 coins There are 16 dimes and 8 quarters total. 16+8=24 Therefore you want to know the probability of drawing one of those 24 coins out of the 50 total coins assuming the size of the coin does not affect the probability that you draw it, then the probability of drawing one of the 24 out of 50 is 24/50 you can simplify this fraction 24/50 = 12 / 25